# TEG-DB: A Comprehensive Dataset and Benchmark of Textual-Edge Graphs

### v1

@@ Abstract
Text-Attributed Graphs (TAGs) augment graph structures with natural language descriptions, facilitating detailed depictions of data and their interconnections across various real-world settings. However, existing TAG datasets predominantly feature textual information only at the nodes, with edges typically represented by mere binary or categorical attributes. This lack of rich textual edge annotations significantly limits the exploration of contextual relationships between entities, hindering deeper insights into graph-structured data. To address this gap, we introduce Textual-Edge Graphs Datasets and Benchmark (TEG-DB), a comprehensive and diverse collection of benchmark textual-edge datasets featuring rich textual descriptions on nodes and edges. The TEG-DB datasets are large-scale and encompass a wide range of domains, from citation networks to social networks. In addition, we conduct extensive benchmark experiments on TEG-DB to assess the extent to which current techniques, including pre-trained language models, graph neural networks, and their combinations, can utilize textual node and edge information. Our goal is to elicit advancements in textual-edge graph research, specifically in developing methodologies that exploit rich textual node and edge descriptions to enhance graph analysis and provide deeper insights into complex real-world networks. The entire TEG-DB project is publicly accessible as an open-source repository on Github, accessible at https://github.com/Zhuofeng-Li/TEG-Benchmark.
**footnotetext: Both authors contributed equally to this work.
@@ 1 Introduction
Text-attributed graphs (TAGs) are graph structures in which nodes are equipped with rich textual information, allowing for deeper analysis and interpretation of complex relationships [47, 18, 16]. TAGs are widely utilized in a variety of real-world applications, including social networks [31, 30], citation networks [25], and recommendation systems [40, 15]. Due to the universal representational capabilities of language, TAGs have emerged as a promising format for potentially unifying a wide range of existing graph datasets. This field has recently garnered rapidly growing interest, particularly in the development of foundational models for graph data [23, 16, 42].
@@
Unfortunately, a central issue in designing the TAG foundation model is the lack of comprehensive datasets with rich textual information on both nodes and edges. Most traditional graph datasets solely offer node attribute embeddings, devoid of the original textual sentences, which results in a significant loss of context and limits the application of advanced techniques such as large language models (LLMs) [24]. Despite some TAG datasets being present recently [42], their data usually only have text information on nodes where the edges are usually represented as binary or categorical. However, the textual information of edges in TAGs is crucial for elucidating the meaning of individual documents and their semantic correlations. For instance, as shown in Figure 1, this scientific article network illustrates the citation patterns of articles authored by Einstein and Planck in the field of quantum mechanics. When we need to conclude that ’Planck endorsed the probabilistic nature of quantum mechanics while Einstein opposed this view,’ and if we consider it in terms of a TAG view, focusing solely on the content of the papers authored by Einstein (Paper A) and Planck (Paper E), we would only conclude that both Einstein and Planck supported quantum mechanics. However, to further deduce that Einstein opposed studying quantum mechanics from a probabilistic perspective, it is necessary to adopt the Textual-Edge Graph (TEG) approach. This approach not only focuses on the paper contents but also pays greater attention to the citation information from the edge between Paper A and Book B, as well as the edge between Paper E and Book D. These citation edges provide essential context and reveal the relationships and influence between different scholarly works.
@@
While compelling, TEGs face three significant challenges that make them an open problem. (1) Comprehensive TEG datasets are absent. Currently, there is a lack of comprehensive TEG datasets that simultaneously incorporate textual information from both nodes and edges, spanning multiple domains of varying sizes, and encompassing various mainstream graph learning tasks. This deficiency hinders the evaluation of TEG-based methods across diverse applications and domains. (2) Existing experimental settings for TEG are disorganized. Due to the inherent variety and complexity of TEGs, coupled with the absence of a standardized data format, existing works have adopted different datasets with different experimental settings [19, 18, 17, 50, 49, 22, 23]. This causes great difficulties in model comparisons in this field. (3) Comprehensive benchmarks and analyses for TEG-based methods are missing. While some techniques can accommodate edge features, they typically process binary or categorical data. It remains unclear if these methods can effectively utilize rich textual information on edges, particularly in leveraging complex interactions between graph nodes.
@@
Present work. Recognizing all the above challenges, our research proposes the Textual-Edge Graphs Datasets and Benchmark (TEG-DB). TEG-DB is a pioneering initiative offering a diverse collection of benchmark graph datasets with rich textual descriptions on both nodes and edges. To address the issue of inadequate TEG datasets, our TEG datasets as shown in Table 1 cover an extensive array of domains, including Book Recommendation, E-commerce, Academic, and Social networks. Ranging in size from small to large, each dataset contains abundant raw text data associated with both nodes and edges, facilitating comprehensive analysis and modeling across various fields. Moreover, to address the inconsistency in experimental settings and the lack of comprehensive analyses for TEG-based methods, we first represent the TEG dataset in a unified format, then conduct extensive benchmark experiments and perform a comprehensive analysis. These experiments are designed to evaluate the capabilities of current computational techniques, such as pre-trained language models and graph neural networks, as well as their integrations. Our contributions are summarized below:
@@
  • We conduct extensive benchmark experiments and perform a comprehensive analysis of TEG-based methods, delving deep into various aspects such as the impact of different models, the effect of embeddings generated by Pre-trained Language Models (PLMs) of various scales, and the influence of different domain datasets. By addressing key challenges and highlighting promising opportunities, our research stimulates and guides future directions for TEG exploration and development.
@@ 2 Related Works
PLM-based methods. PLM-based methods leverage the power of PLM to enhance the text modeling within each node due to their pre-training on a vast corpus. The early works on modeling textual attributes were based on shallow networks, e.g., Skip-Gram [28] and GloVe [32]. In recent years, Large Language Models (LLM) have become trending tools. Models like Llama [36], PaLM [3], and GPT [1] show their strong comprehension and inferring ability in cross-field natural language based tasks like code generation [4], legal consulting [7], make creative arts [21], as well as understanding and learning from Graphs [6]. One of the key applications of pre-trained language models is text representation, in which low-dimensional embeddings capture the underlying semantics of texts. On the TAGs, the PLMs use the local textual information of each node to learn a good representation for the downstream task.
@@
GNN-based methods. The rapid advancements in graph representation learning within machine learning have led to numerous studies addressing various tasks, such as node classification [20] and link prediction [48]. Graph neural networks (GNNs) are acknowledged as robust tools for modeling graph data. These methods, including GCN [20], GAT [37], GraphSAGE [11], GIN [41], and RevGAT [29], develop effective message-passing mechanisms that facilitate information aggregation between nodes, thereby enhancing graph representations. GNNs typically utilize the "cascade architecture" advocated by GraphSAGE for textual graph representation, wherein node features are initially encoded independently using text modeling tools (e.g., PLMs) and then aggregated by GNNs to generate the final representation.
@@
LLM as Predictor. In recent years, several recent studies [43, 5, 10] have delved into the potential of Large Language Models (LLMs) in analyzing graph-structured data. However, there is a lack of comprehensive research on the ability of LLMs to effectively identify and utilize key topological structures across various prompt scenarios, task complexities, and datasets. Chen et al. [5] and Guo et al. [10] proposed using LLMs on graph data but primarily focused on node classification within specific citation network datasets, limiting the exploration of LLMs’ performance across various tasks and datasets. Furthermore, Ye et al. [43] fine-tuned LLMs on a specific dataset to outperform GNNs, focusing on a different research goal, which emphasizes LLMs’ inherent ability to understand and leverage graph structures.
Benchmarks for text-attributed graphs. We conducted an extensive survey of various text-attributed graph datasets previously utilized in the literature. Our observations reveal that the majority of traditional node-level datasets are essentially text-attributed graphs. However, these datasets have shortcomings in representation learning on TAGs due to a lack of raw textual information. Recently, certain TAG datasets [42, 19] have emerged to address these limitations. However, they still have shortcomings when exploring representation learning on TEGs. Firstly, these datasets typically only include text information on nodes, while edges are often represented as binary or categorical, limiting a comprehensive understanding of node relationships. Secondly, they lack coverage across diverse domains and tasks, potentially hindering the discovery of robust and generalizable models. Lastly, their representation formats lack uniformity, introducing inconsistencies and complexities in analysis and modeling techniques across datasets.
@@ 3 Preliminaries
Dataset Nodes Edges Nodes-Class Graph Domain Size Nodes-text Edges-text Node Classification Link Prediction Previous Twitch Social Network [33] 7,126 88,617 2 Social Networks Small Facebook Page-Page Network [34] 22,470 171,002 4 Social Networks Small ogbn-arxiv [13] 169,343 1,166,243 40 Academic Medium Citeseer [35] 3,327 4,732 6 Academic Small Pubmed [35] 19,717 44,338 3 Academic Small Cora [27] 2,708 5,429 7 Academic Small CitationV8 [42] 1,106,759 6,120,897 - Academic Large GoodReads [42] 676,084 8,582,324 11 Book Recommendation Large Sports-Fitness [42] 173,055 1,773,500 13 E-commerce Medium Ele-Photo [42] 48,362 500,928 12 E-commerce Small Books-History [42] 41,551 358,574 12 E-commerce Small Books-Children [42] 76,875 1,554,578 24 E-commerce Small ogbn-arxiv-TA [42] 169,343 1,166,243 40 Academic Medium Ours Goodreads-History 540,807 2,368,539 11 Book Recommendation Large Goodreads-Crime 422,653 2,068,223 11 Book Recommendation Large Goodreads-Children 216,624 858,586 11 Book Recommendation Large Goodreads-Comics 148,669 631,649 11 Book Recommendation Medium Amazon-Movie 137,411 2,724,028 399 E-commerce Medium Amazon-Apps 31,949 62,036 62 E-commerce Small Reddit 478,022 676,684 3 Social Networks Large Twitter 18,761 23,764 503 Social Networks Small Citation 4,972,456 5,970,965 24 Academic Large
@@ 4.2 Data Preparation and Construction
In order to construct the dataset with simultaneous satisfaction of both rich textual information on nodes and edges, nine datasets from diverse domains and different scales are chosen. Specifically, we collect four User-Book Review networks from Goodreads datasets [38] in the Book Recommendation domain and two shopping networks from Amazon datasets [12] in the E-commerce domain. One citation network from the Open Research Corpus [2] in the academic domain. Two social networks from Reddit [11] and Twitter [46]. The statistics of the datasets are shown in Table 1.
@@
The creation of text-attributed graph datasets involves three main steps. Firstly, preprocessing the textual attributes within the original dataset, which includes tasks such as handling missing values, filtering out non-English statements, removing anomalous symbols, and truncating excessive length. Secondly, constructing the graph itself. The connectivity between nodes is derived from inherent relationships provided within the dataset, such as citation relationships between papers in citation networks. It is important to note that during graph construction, self-edges and isolated nodes are eliminated. Lastly, refining the constructed graph. It is noteworthy that our dataset encompasses all major tasks in graph representation learning: node classification and link prediction. Below are the specifics of each dataset:
@@
User-Book Review Network. Four datasets within the realm of User-Book Review Networks, specifically labeled as Goodreads-History, Goodreads-Crime, Goodreads-Children, and Goodreads-Cosmics, were formulated. The Goodreads datasets [38] are the main source. Nodes represent different types of books and reviewers, while edges indicate book reviews. Node labels are assigned based on the book category. The descriptions of books are used as node textual information, and reviews of users are used as edges textual information. The corresponding tasks are to predict the categories of the books, which is formulated as a multi-label classification problem, and to predict whether there are connections between users and books. These comprehensive data help infer user preferences and identify similar tastes, enhancing online book recommendations, unlike existing datasets that often lack interaction texts.
@@
Shopping Networks. Two datasets, Amazon-Apps and Amazon-Movie, are classified under Shopping Networks. The Amazon item dataset [12] is the primary source, encompassing item reviews and descriptions. Nodes represent different types of items and reviewers, while edges indicate item reviews. The descriptions of items are used as node textual information, and reviews of users are used as edge textual information. The corresponding tasks are to predict the categories of the items, formulated as a multi-label classification problem, and to predict whether there are connections between users and items. These datasets have the potential to significantly enhance recommendation systems, providing richer data for more accurate suggestions and a personalized shopping experience.
@@
Citation Networks. The raw data for the citation network is sourced from the Open Research Corpus, derived from the complete Semantic Scholar corpus [2]. Nodes represent papers and edges represent the citation relationships between papers. The descriptions of papers are used as node textual information, and citation information, such as the context and paragraphs in which papers are cited, is utilized as textual edge data. The corresponding task involves predicting the domain to which a paper belongs, formulated as a multi-class classification problem, and predicting there exists a citation relationship between papers. This dataset enhances academic network expressiveness, particularly benefiting tasks like node classification and link prediction in graph machine learning.
@@
Social Networks. The Reddit dataset, sourced from Reddit [11], and the Twitter dataset, derived from Twitter [46], represent two prominent social media platforms. Nodes represent users and topics. The edges in the dataset indicate two types of reviews: those between users, representing user-user links, and those between users and topics, representing user-topic links. The descriptions of topics are used as node textual information, and reviews are used as edge textual information. The corresponding tasks are to predict the categories of the topics, formulated as a multi-class classification problem, and to predict whether there are connections between users and topics. These datasets can enhance recommendation systems and offer a more personalized shopping experience. Utilizing these datasets enhances recommendation algorithm performance, providing more personalized and relevant suggestions, while also offering valuable insights into user interests and preferences for social network research and business decision-making.

### v3

@@ Abstract
Text-Attributed Graphs (TAGs) augment graph structures with natural language descriptions, facilitating detailed depictions of data and their interconnections across various real-world settings. However, existing TAG datasets predominantly feature textual information only at the nodes, with edges typically represented by mere binary or categorical attributes. This lack of rich textual edge annotations significantly limits the exploration of contextual relationships between entities, hindering deeper insights into graph-structured data. To address this gap, we introduce Textual-Edge Graphs Datasets and Benchmark (TEG-DB), a comprehensive and diverse collection of benchmark textual-edge datasets featuring rich textual descriptions on nodes and edges. The TEG-DB datasets are large-scale and encompass a wide range of domains, from citation networks to social networks. In addition, we conduct extensive benchmark experiments on TEG-DB to assess the extent to which current techniques, including pre-trained language models (PLMs), graph neural networks (GNNs), proposed novel entangled GNNs and their combinations, can utilize textual node and edge information. Our goal is to elicit advancements in textual-edge graph research, specifically in developing methodologies that exploit rich textual node and edge descriptions to enhance graph analysis and provide deeper insights into complex real-world networks. The entire TEG-DB project is publicly accessible as an open-source repository on Github, accessible at https://github.com/Zhuofeng-Li/TEG-Benchmark.
@@ 1 Introduction
Text-attributed graphs (TAGs) are graph structures in which nodes are equipped with rich textual information, allowing for deeper analysis and interpretation of complex relationships Zhang et al. (2024); Jin et al. (2023c, a). TAGs are widely utilized in a variety of real-world applications, including social networks Paranyushkin (2019); Myers et al. (2014), citation networks Liu et al. (2013), and recommendation systems Wu et al. (2022); Huang et al. (2004). Due to the universal representational capabilities of language, TAGs have emerged as a promising format for potentially unifying a wide range of existing graph datasets. This field has recently garnered rapidly growing interest, particularly in the development of foundational models for graph data Ling et al. (2024); Jin et al. (2023a); Yan et al. (2023).
@@
Unfortunately, a central issue in designing the TAG foundation model is the lack of comprehensive datasets with rich textual information on both nodes and edges. Most traditional graph datasets solely offer node attribute embeddings, devoid of the original textual sentences, which results in a significant loss of context and limits the application of advanced techniques such as large language models (LLMs) Ling et al. (2023). Despite some TAG datasets being present recently Yan et al. (2023), their data usually only have text information on nodes where the edges are usually represented as binary or categorical. However, the textual information of edges in TAGs is crucial for elucidating the meaning of individual documents and their semantic correlations. For instance, as shown in Figure 1, this scientific article network illustrates the citation patterns of articles authored by Einstein and Planck in the field of quantum mechanics. When we need to conclude that ’Planck endorsed the probabilistic nature of quantum mechanics while Einstein opposed this view,’ and if we consider it in terms of a TAG view, focusing solely on the content of the papers authored by Einstein (Paper A) and Planck (Paper E), we would only conclude that both Einstein and Planck supported quantum mechanics. However, to further deduce that Einstein opposed studying quantum mechanics from a probabilistic perspective, it is necessary to adopt the Textual-Edge Graph (TEG) approach. This approach not only focuses on the paper contents but also pays greater attention to the citation information from the edge between Paper A and Book B, as well as the edge between Paper E and Book D. These edges provide essential citation context and reveal the relationships and influence between different scholarly works.
@@
While compelling, TEGs face three significant challenges that make them an open problem. (1) Comprehensive TEG datasets are absent. Currently, there is a lack of comprehensive TEG datasets that simultaneously incorporate textual information from both nodes and edges, spanning multiple domains of varying sizes, and encompassing various mainstream graph learning tasks. This deficiency hinders the evaluation of TEG-based methods across diverse applications and domains. (2) Existing experimental settings for TEG are disorganized. Due to the inherent variety and complexity of TEGs, coupled with the absence of a standardized data format, existing works have adopted different datasets with different experimental settings Jin et al. (2023d, c, b); Zou et al. (2023); Zhao et al. (2022); Kuang et al. (2023); Ling et al. (2024). This causes great difficulties in model comparisons in this field. (3) Comprehensive benchmarks and analyses for TEG-based methods are missing. While some techniques can accommodate edge features, they typically process binary or categorical data. It remains unclear if these methods can effectively utilize rich textual information on edges, particularly in leveraging complex interactions between graph nodes.
@@
Present work. Recognizing all the above challenges, our research proposes the Textual-Edge Graphs Datasets and Benchmark (TEG-DB). TEG-DB is a pioneering initiative offering a diverse collection of benchmark graph datasets with rich textual descriptions on both nodes and edges. To address the issue of inadequate TEG datasets, our TEG datasets as shown in Table 1 cover an extensive array of domains, including Book Recommendation, E-commerce, Academic, and Social networks. Ranging in size from small to large, each dataset contains abundant raw text data associated with both nodes and edges, facilitating comprehensive analysis and modeling across various fields. Moreover, to address the inconsistency in experimental settings and the lack of comprehensive analyses for TEG-based methods, we first represent the TEG dataset in a unified format, then conduct extensive benchmark experiments and perform a comprehensive analysis. These experiments are designed to evaluate the capabilities of current computational techniques, such as pre-trained language models (PLMs), graph neural networks (GNNs) and proposed novel entangled GNNs, as well as their integrations. Our contributions are summarized below:
@@
  • We conduct extensive benchmark experiments and perform a comprehensive analysis of TEG-based methods, delving deep into various aspects such as the impact of different models and embeddings generated by PLMs of various scales, the consequence of diverse embedding methods in GNNs including separate and entangled embeddings, the effect of edge text and the influence of different domain datasets. By addressing key challenges and highlighting promising opportunities, our research stimulates and guides future directions for TEG exploration and development.
@@ 2 Related Works
PLM-based methods. PLM-based methods leverage the power of PLM to enhance the text modeling within each node due to their pre-training on a vast corpus. The early works on modeling textual attributes were based on shallow networks, e.g., Skip-Gram Mikolov et al. (2013) and GloVe Pennington et al. (2014). In recent years, Large Language Models (LLM) have become trending tools. Models like Llama Touvron et al. (2023), PaLM Anil et al. (2023), and GPT Achiam et al. (2023) show their strong comprehension and inferring ability in cross-field natural language based tasks like code generation Cai et al. (2023), legal consulting Cui et al. (2023), make creative arts Koh et al. (2023), as well as understanding and learning from Graphs Chen et al. (2024). One of the key applications of pre-trained language models is text representation, in which low-dimensional embeddings capture the underlying semantics of texts. On the TAGs, the PLMs use the local textual information of each node to learn a good representation for the downstream task.
@@
GNN-based methods. The rapid advancements in graph representation learning within machine learning have led to numerous studies addressing various tasks, such as node classification Kipf and Welling (2017) and link prediction Zhang and Chen (2018). Graph neural networks (GNNs) are acknowledged as robust tools for modeling graph data. These methods, including GCN Kipf and Welling (2017), GAT Velickovic et al. (2018), GraphSAGE Hamilton et al. (2017), GIN Xu et al. (2019), and RevGAT Ming Chen (2020), develop effective message-passing mechanisms that facilitate information aggregation between nodes, thereby enhancing graph representations. GNNs typically utilize the "cascade architecture" advocated by GraphSAGE for textual graph representation, wherein node features are initially encoded independently using text modeling tools (e.g., PLMs) and then aggregated by GNNs to generate the final representation.
@@
LLM as Predictor. In recent years, several recent studies Ye et al. (2023); Chen et al. (2023); Guo et al. (2023) have delved into the potential of Large Language Models (LLMs) in analyzing graph-structured data. However, there is a lack of comprehensive research on the ability of LLMs to effectively identify and utilize key topological structures across various prompt scenarios, task complexities, and datasets. Chen et al. Chen et al. (2023) and Guo et al. Guo et al. (2023) proposed using LLMs on graph data but primarily focused on node classification within specific citation network datasets, limiting the exploration of LLMs’ performance across various tasks and datasets. Furthermore, Ye et al. Ye et al. (2023) fine-tuned LLMs on a specific dataset to outperform GNNs, focusing on a different research goal, which emphasizes LLMs’ inherent ability to understand and leverage graph structures.
Benchmarks for text-attributed graphs. Current benchmarks in text-attributed graph representation learning can be divided into two stages. The first stage benchmark includes datasets such as mag Wang et al. (2020) and ogbn-arxiv Hu et al. (2020a), which feature limited textual information primarily associated with nodes. The second stage benchmark is represented by CS-TAG Yan et al. (2023), which builds upon the first stage by providing richer node-level textual data. However, these datasets face limitations in exploring representation learning for textual-edge graphs. Specifically, they typically include text only on nodes, with edges often represented as binary or categorical, which restricts a comprehensive understanding of node semantic relationships. Additionally, they lack coverage across diverse domains and tasks, hindering the development of robust and generalizable models. Furthermore, the lack of uniformity in representation formats introduces inconsistencies and complexities in analysis and modeling. Thus, there is a clear need for the development of a comprehensive benchmark with textual information on both nodes and edges in a unified format.
@@ 3 Preliminaries
Dataset Nodes Edges Nodes-Class Graph Domain Size Nodes-text Edges-text Node Classification Link Prediction Previous Twitch Social Network Rozemberczki et al. (2019) 7,126 88,617 2 Social Networks Small Facebook Page-Page Network Rozemberczki et al. (2020) 22,470 171,002 4 Social Networks Small ogbn-arxiv Hu et al. (2020a) 169,343 1,166,243 40 Academic Medium Citeseer Sen et al. (2008) 3,327 4,732 6 Academic Small Pubmed Sen et al. (2008) 19,717 44,338 3 Academic Small Cora McCallum et al. (2000) 2,708 5,429 7 Academic Small CitationV8 Yan et al. (2023) 1,106,759 6,120,897 - Academic Large GoodReads Yan et al. (2023) 676,084 8,582,324 11 Book Recommendation Large Sports-Fitness Yan et al. (2023) 173,055 1,773,500 13 E-commerce Medium Ele-Photo Yan et al. (2023) 48,362 500,928 12 E-commerce Small Books-History Yan et al. (2023) 41,551 358,574 12 E-commerce Small Books-Children Yan et al. (2023) 76,875 1,554,578 24 E-commerce Small ogbn-arxiv-TA Yan et al. (2023) 169,343 1,166,243 40 Academic Medium Ours Goodreads-History 540,807 2,368,539 11 Book Recommendation Large Goodreads-Crime 422,653 2,068,223 11 Book Recommendation Large Goodreads-Children 216,624 858,586 11 Book Recommendation Large Goodreads-Comics 148,669 631,649 11 Book Recommendation Medium Amazon-Movie 137,411 2,724,028 399 E-commerce Medium Amazon-Apps 31,949 62,036 62 E-commerce Small Reddit 478,022 676,684 3 Social Networks Large Twitter 18,761 23,764 505 Social Networks Small Citation 169,343 1,166,243 40 Academic Large
@@ 4.2 Data Preparation and Construction
In order to construct the dataset with simultaneous satisfaction of both rich textual information on nodes and edges, nine datasets from diverse domains and different scales are chosen. Specifically, we collect four User-Book Review networks from Goodreads datasets Wan and McAuley (2018); Wan et al. (2019) in the Book Recommendation domain and two shopping networks from Amazon datasets He and McAuley (2016); Hou et al. (2024) in the E-commerce domain. Two social networks from Reddit and Twitter McMinn et al. (2013). One citation network from MAG Wang et al. (2020) and The Semantic Scholar Open Data Platform Kinney et al. (2023) in the academic domain. The statistics of the datasets are shown in Table 1.
@@
The creation of textual-edge graph datasets involves three main steps. Firstly, preprocessing the textual attributes within the original dataset, which includes tasks such as handling missing values, filtering out non-English statements, removing anomalous symbols, truncating excessive length and selecting the most relevant textual attributes as the raw text for nodes or edges. Secondly, constructing the TEG itself. The connectivity between nodes is derived from inherent relationships provided within the dataset, such as citation relationships between papers in citation networks. It is important to note that during graph construction, self-edges and isolated nodes are eliminated. Lastly, refining the constructed graph. It is noteworthy that our dataset encompasses all major tasks in graph representation learning: node classification and link prediction. Below are the specifics of each dataset:
@@
User-Book Review Network. Four datasets within the realm of User-Book Review Networks, specifically labeled as Goodreads-History, Goodreads-Crime, Goodreads-Children, and Goodreads-Cosmics, were formulated. The Goodreads datasets are the main source. Nodes represent different types of books and reviewers, while edges indicate book reviews. Node labels are assigned based on the book categories. The descriptions of books are used as book node textual information while user information serves as the user node textual information and reviews of users are used as edges textual information. The corresponding tasks are to predict the categories of the books, which is formulated as a multi-label classification problem, and to predict whether there are connections between users and books. These comprehensive data help infer user preferences and identify similar tastes, enhancing online book recommendations, unlike existing datasets that often lack interaction texts.
@@
Shopping Networks. Two datasets, Amazon-Apps and Amazon-Movie, are classified under Shopping Networks. The Amazon datasets are the primary source, encompassing item reviews and descriptions. Nodes represent different types of items and reviewers, while edges indicate item reviews. The descriptions of items are used as item node textual information, while user information serves as the user node textual information and reviews of users are used as edge textual information. The corresponding tasks are to predict the categories of the items, formulated as a multi-label classification problem, and to predict whether there are connections between users and items. These datasets have the potential to significantly enhance recommendation systems, providing richer data for more accurate suggestions and a personalized shopping experience.
@@
Citation Networks. The raw data for the citation network is sourced from the MAG and The Semantic Scholar Open Data Platform. Nodes represent papers, and edges represent the citation relationship. The titles and abstracts of papers are used as node textual information, and citation information, such as the context and paragraphs in which papers are cited, is utilized as textual edge data. The corresponding task involves predicting the domain to which a paper belongs, formulated as a multi-class classification problem, and predicting whether there exists a citation relationship between papers. This dataset enhances academic network expressiveness, particularly benefiting tasks like node classification and link prediction in graph machine learning.
@@
Social Networks. The Reddit dataset, sourced from Reddit and the Twitter dataset, derived from Twitter, represent two prominent social media platforms. Nodes represent users and topics. The edges indicate the post-relationship. The descriptions of topics are used as topic node textual information while user information serves as the user node textual information and post text in subreddits or tweets is used as edge textual information. The corresponding tasks are to predict the category of the topics, formulated as a multi-class classification problem, and to predict whether there are connections between users and topics. Utilizing these datasets enhances recommendation algorithm performance, providing more personalized and relevant suggestions, while also offering valuable insights into user interests and preferences for social network research and business decision-making.